OpenStudy (anonymous):

A system of equations is shown below: 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

for Part A would i do elimination? 6x+16y=24 -6x-6y=-9 10y=15 y=1.5?

OpenStudy (aum):

You need to solve the equation only in Part B.

OpenStudy (aum):

For Part A: "replacing one equation with the sum of that equation and a multiple of the other" replace first equation by sum of the first equation and some constant times the second equation. After that, replace second equation by sum of the second equation and some constant times the first equation.

OpenStudy (aum):

You can choose the constant as 2. Replace (3x + 8y = 12) by (3x + 8y = 12) + 2 * (2x + 2y = 3) Simplify the above.

OpenStudy (anonymous):

(3x+8y=12)+(4x+4y=6) 7x+12y=18?

OpenStudy (aum):

Correct.

OpenStudy (anonymous):

so that will be the equation for part A?

OpenStudy (aum):

You have to do a similar operation with the second equation. replace second equation by sum of the second equation and some constant times the first equation.

OpenStudy (aum):

Replace (2x + 2y = 3) by (2x + 2y = 3) + 2 * (3x + 8y = 12)

OpenStudy (anonymous):

(2x+2y=3)+(6x+16y=24) 8x+18y=27

OpenStudy (aum):

Part A) The equivalent system of equations are: 7x +12y =18 8x +18y = 27

OpenStudy (anonymous):

so the equivalent systems of equations will be 8x+18y=27 7x+12y=18 Now find the value for x and y?

OpenStudy (aum):

Part B) First solve the original system of equations. Then solve the equivalent system of equations created in part A. Prove the solutions for x and y are the same in both cases.

OpenStudy (aum):

Original system of equations: 3x + 8y = 12 2x + 2y = 3 Multiply second equation by -4 and add it to the first. Solve for y.

OpenStudy (aum):

solve for x. (i meant)

OpenStudy (anonymous):

i got x=0?

OpenStudy (aum):

Yes. Sub x = 0 in the second equation and find y.

OpenStudy (anonymous):

oh lol thought i made a mistake y=1.5

OpenStudy (anonymous):

(0,1.5) for original system

OpenStudy (aum):

Correct. Or in fractions: (0, 3/2). Solve the equivalent system created in part A.

OpenStudy (anonymous):

8x+12y=18 8x+18y=27

OpenStudy (anonymous):

no 7x+12y=18 8x+18y=27

OpenStudy (anonymous):

-54x-96y=-144 54x+144y=216 48y=72 y=1.5

OpenStudy (anonymous):

7x+12(1.5)=18 7x+18=18 7x=0 x=0

OpenStudy (aum):

you got the correct y, but 7 * 8 = 56 (not 54)

OpenStudy (anonymous):

oops my mistake

OpenStudy (aum):

-56x-96y=-144 56x+144y=216 48y=72 y=1.5

OpenStudy (anonymous):

yes thats what i meant hit 4 on accident and copied for the second equation

OpenStudy (anonymous):

thank you

OpenStudy (aum):

you are welcome.